OpenStudy (anonymous):

Solve the system of linear equations by the substitution method x = 3y + 30 3x + 6y = 0 Can someone show me the steps as to how to do this? Thanks!

OpenStudy (aonz):

x = 3y + 30 3x + 6y = 0 now we sub x into the 2nd equation since x = 3y + 30 3(3y + 30) = 0

OpenStudy (aonz):

3(3y +30 ) +6y = 0 sorry

OpenStudy (aonz):

9y + 90 + 6y = 0 15y + 90 = 0 15y = -90 y =-6 now can you find x value?

OpenStudy (espex):

Good job demonstrating and not just handing out the answer, @AonZ.